# dUltra: Ultra-Fast Diffusion Language Models via Reinforcement Learning

### v1

@@ 4.2.2 Results with Large Block Size
Fast-dLLM 75.66 54.43 32.60 75.12 35.40 61.54 37.80 56.93
dParallel 75.66 24.07 31.40 28.49 30.49 23.68 36.00 18.07
d3LLM 69.07 21.37 31.40 26.65 32.32 24.49 33.80 26.37
dUltra-math-b128 (ours) 77.65 16.38 33.64 26.53 - - - -
   w/o planner 69.67 10.46 32.20 19.36 - - - -
dUltra-coding-b128 (ours) 79.48 20.15 34.20 30.90 36.83 21.79 39.12 14.78
   w/o planner 71.57 10.90 30.60 18.46 21.95 13.18 21.60 8.27
Table 2: Main results comparing dUltra with baseline methods across mathematical reasoning and code generation tasks. The generation length is 256 and the block size is 128 for all results in this table. We report accuracy and NFE (number of function evaluations) for each task. Bold indicates the best result and indicates the second-best result for each column. The “w/o planner” rows show results using Fast-dLLM sampling with the base model’s logits (without the learned unmasking planner head), serving as an ablation baseline. dUltra-math-b128 is trained on GSM8K training dataset with rollouts using block size of 128, and uses dUltra-math as the base checkpoint. dUltra-coding-b128 is trained on the APPS dataset with rollouts using block size of 128, and uses dUltra-coding as the base checkpoint.
@@
4.3 Efficiency-Performance Trade-off
The unmasking probability predicted by the unmasking planner head directly controls the expected number of tokens unmasked at each denoising step. This allows us to strike a balance between sampling efficiency (NFE) and task performance (accuracy). By multiplying the unmasking probability by a factor $\alpha$, we are able to adjust the sampling speed at inference time without retraining the model. Figure˜6 illustrates the accuracy vs. NFE trade-off on all tasks for both block size 32 and block size 128.
@@ 4.4 Analysis of Learned Unmasking Strategies
Figure 7: Visualization of learned unmasking patterns on GSM8K. Red bars show newly unmasked tokens at each denoising step. The pattern exhibits a predominantly autoregressive (left-to-right) trend with minor deviations for tokens with strong contextual support.
@@ 5 Related Work
Reinforcement Learning for DLMs. Recent work has explored post-training diffusion language models (dLLMs) with reinforcement learning to improve their reasoning and generation capabilities. Early examples include d1, which adapts GRPO-style optimization to the (approximate) likelihood structure of dLLMs (zhao2025d1), and DiffuCoder, which targets code generation for LLaDA-style masked diffusion models (gong2025diffucoder). A growing line of follow-up work focuses on improving the stability and bias/variance trade-offs of policy-gradient estimators for dLLMs (tang_wd1_2025; rojas2025improving; lin2025boundary; wang2025spg; zhu2025enhancing; mmada). Notably, wang2025revolutionizing; zhao2025diffpo try to obtain more accurate likelihood by using more intermediate decoding steps. While these methods focus on reasoning capabilities, we propose to learn the optimal unmasking order and enable high quality mode filtering via RL. Close to our work is DCOLT (huang2025dcolt), which also learns an unmasking order policy with RL. However, their unmasking policy does not enable parallel decoding due to the adoption of the Plackett-Luce model. zhao2025diffpo learns adaptive confidence thresholds for unmasking with RL, but their method still relies on logit-based heuristics for token selection rather than learning a flexible unmasking policy as in our work. Concurrent work jazbec2025 also learns an unmasking policy head with GRPO, but there are several important distinctions. First of all, they learn the unmasking policy from the logits rather than the last layer representation. Secondly, they do not jointly train the base model and the unmasking policy, but rather keep the base model fixed. Finally, they do not introduce an on-policy distillation reward to encourage mode-seeking behavior. As a result, they report only marginal benefits over logit-based heuristics, while we achieve better performance in both accuracy and NFE even when compared to SOTA distillation-based methods.
@@
DLM Acceleration. Most existing acceleration methods for masked dLLMs are either system level, such as enabling kv-cache and employing token dropping (ma2025dkv; liu2025dllm; hu2025acceleratingdiffusionlanguagemodel; wu2025fastdllmtrainingfreeaccelerationdiffusion) or algorithmic level that designs better sampling strategies. Among the latter, there are two approaches: (i) training-free logit-based heuristic decoders (wu2025fastdllmtrainingfreeaccelerationdiffusion; israel2025acceleratingdiffusionllmsadaptive; wei2025acceleratingdiffusionlargelanguage; ben-hamu_accelerated_2025) or (ii) distillation-based approaches that fit the off-policy trajectories produced by a base model (chen2025dparallellearnableparalleldecoding; d3llm). Heuristic samplers often rely on confidence, entropy or other related uncertainty signals computed from the logits, and have recently enabled substantial inference speedups in dLLMs via parallel decoding and KV caching (wu2025fastdllmtrainingfreeaccelerationdiffusion; hu2025acceleratingdiffusionlanguagemodel). Distillation-based methods, such as pseudo-trajectory distillation in d3LLM (d3llm) and learnable parallel decoding in dParallel (chen2025dparallellearnableparalleldecoding), can further increase decoding efficiency by finetuning the base model to sample multiple steps of the base model in one step, which essentially filters the modes of the base model. These methods depend on the quality and coverage of the teacher trajectories, which also may become off-policy as training progresses. Our approach explicitly addresses this mismatch by using on-policy rollouts with GRPO, while also encouraging mode seeking behavior.
@@
Denoising Trajectory Planning. Beyond heuristic confidence-based planning, recent works have explored learned planning modules for MDLM trajectories. These approaches differ fundamentally in their planning strategies. P2 (peng2025pathplanningmaskeddiffusion) introduces explicit planning and denoising sub-stages, enabling iterative refinement of existing unmasked tokens. DCOLT (huang2025dcolt) employs the Plackett-Luce model to generate sequential unmasking orders optimized for accuracy rather than parallel decoding efficiency. DiFFPO (zhao2025diffpo) learns adaptive per-prompt thresholds via off-policy RL but still selects tokens based on confidence heuristics. In the concurrent work, jazbec2025 learns an unmasking policy head with GRPO on token confidences from logits.
@@ 6 Conclusion
We present dUltra, a framework for learning optimal unmasking strategies to accelerate masked diffusion language models through reinforcement learning. Motivated by the need for mode filtering and maximizing conditionally independent tokens per step, we train an unmasking planner head jointly with the base model using GRPO with on-policy rollouts and distillation rewards. Empirically, dUltra outperforms distillation baselines (d3LLM, dParallel) in both accuracy and NFEs, validating the superiority of our method. We found that learned unmasking order adapts to the task structure naturally: code generation exploits more parallelism locally by learning an unmasking order resembling the random masking strategy while math reasoning decodes multiple consecutive tokens sequential and in parallel, exhibiting mode filtering behaviors. This work demonstrates that learned unmasking policies guided by mode filtering and conditional independence principles enable progress towards "diffusion supremacy" over autoregressive LLMs.
@@ A.1.1 Model Architecture
Our model extends the LLaDA-Instruct model [nie2025largelanguagediffusionmodels] with an unmasking planner head. The architecture consists of:
Base Model. We use LLaDA [nie2025largelanguagediffusionmodels], a masked diffusion language model with Llama-like architecture adapted for discrete diffusion. The model uses a special mask token and employs time-agnostic parametrization following recent findings that the optimal ELBO solution is independent of the time variable [ou2025absorbingdiscretediffusionsecretly, sahoo_simple_2024].
Unmasking Planner Head. We introduce an unmasking planner head that takes the last-layer hidden states and outputs per-token unmasking probabilities:

### v2

@@ 4.2.2 Results with Large Block Size
Fast-dLLM 75.66 54.43 (4.56) 32.60 75.72 (3.30) 35.40 61.54 (4.08) 37.80 56.93 (4.07)
dParallel 75.66 24.07 (8.23) 31.40 28.49 (5.45) 30.49 23.68 (6.26) 36.00 18.07 (7.63)
d3LLM 69.07 21.37 (9.59) 31.40 26.65 (7.03) 32.32 24.49 (7.34) 33.80 26.37 (7.66)
dUltra-math-b128 (ours) 77.65 16.38 (13.96) 33.64 26.53 (9.25) - - - -
   w/o planner 69.67 10.46 (22.16) 32.80 19.36 (12.80) - - - -
dUltra-coding-b128 (ours) 79.48 20.15 (9.99) 34.20 30.90 (7.50) 36.83 21.79 (9.43) 39.12 14.78 (9.97)
   w/o planner 71.57 10.90 (18.44) 30.60 18.46 (12.50) 21.95 13.18 (15.00) 21.60 8.27 (17.65)
Table 2: Main results comparing dUltra with baseline methods across mathematical reasoning and code generation tasks. The generation length is 256 and the block size is 128 for all results in this table. We report accuracy, NFE (number of function evaluations), and T/F (tokens per forward pass) for each task. Bold indicates the best result and indicates the second-best result for each column. The “w/o planner” rows show results using Fast-dLLM sampling with the base model’s logits (without the learned unmasking planner head), serving as an ablation baseline. dUltra-math-b128 is trained on GSM8K training dataset with rollouts using block size of 128, and uses dUltra-math as the base checkpoint. dUltra-coding-b128 is trained on the APPS dataset with rollouts using block size of 128, and uses dUltra-coding as the base checkpoint.
@@
4.3 Speed-Quality Trade-off
During inference time, to eliminate randomness and dynamically adjust the sampling speed, we use a deterministic threshold to only unmask tokens with unmasking probability above that threshold. This allows us to strike a balance between sampling efficiency (NFE) and task performance (accuracy). We use Fast-dLLM’s factor-based parallel decoding strategy (Wu et al., 2025) for models without a planner. We sweep across the same set of thresholds (for dUltra models) and factors (for models w/o a planner) and show the Pareto frontier in Figure˜6. We also show that continual training of dUltra-math on coding tasks does not alter the Pareto curve significantly (see Figure˜10). Furthermore, we explore another trade-off strategy with a multiplicative factor applied to the unmasking probability in Appendix C.
Figure 6: Accuracy-efficiency Pareto frontiers across math reasoning and code generation benchmarks. We compare variants of dUltra against baseline methods on GSM8K (top-left), MATH500 (top-right), HumanEval (bottom-left), and MBPP (bottom-right). The x-axis shows average NFE (lower is faster), while the y-axis shows task accuracy. Each point represents a different threshold for dUltra models with planners and a different factor for baseline models without planners (including dUltra models without planners). dUltra models with planner consistently achieves superior accuracy-efficiency trade-offs compared to distillation-based methods, demonstrating the effectiveness of our learned unmasking strategies via on-policy distillation.
@@ 4.4 Analysis of Learned Unmasking Strategies
Figure 7: Visualization of learned unmasking patterns on GSM8K. Red bars show newly unmasked tokens at each denoising step. The pattern exhibits a predominantly autoregressive (left-to-right) trend with minor deviations. An interesting observation is that sometimes, "Step" and ":" are generated first without the numbering in between, and the numbering is generated in the next step.
@@ 5 Related Work
Reinforcement Learning for DLMs. Recent work has explored post-training diffusion language models (dLLMs) with reinforcement learning to improve their reasoning and generation capabilities. Early examples include d1, which adapts GRPO-style optimization to the (approximate) likelihood structure of dLLMs (Zhao et al., 2025b), and DiffuCoder, which targets code generation for LLaDA-style masked diffusion models (Gong et al., 2025). A growing line of follow-up work focuses on improving the stability and bias/variance trade-offs of policy-gradient estimators for dLLMs (Tang et al., 2025; Rojas et al., 2025; Lin et al., 2025; Wang et al., 2025a; Zhu et al., 2025; Yang et al., 2025). Notably, Wang et al. (2025b); Zhao et al. (2025a) try to obtain more accurate likelihood by using more intermediate decoding steps. While these methods focus on reasoning capabilities, we propose to learn the optimal unmasking order and enable high quality mode filtering via RL. Close to our work is DCOLT (Huang et al., 2025), which also learns an unmasking order policy with RL. However, their unmasking policy does not enable parallel decoding due to the adoption of the Plackett-Luce model. Zhao et al. (2025a) learns adaptive confidence thresholds for unmasking with RL, but their method still relies on logit-based heuristics for token selection rather than learning a flexible unmasking policy as in our work. Concurrent work Jazbec et al. (2025) also learns an unmasking policy head with GRPO, but there are several important distinctions. First of all, they learn the unmasking policy from the logits rather than the last layer representation. Secondly, they do not jointly train the base model and the unmasking policy, but rather keep the base model fixed. Finally, they do not introduce an on-policy distillation reward to encourage mode-seeking behavior. As a result, they report only marginal benefits over logit-based heuristics, while we achieve better performance in both accuracy and NFE even when compared to SOTA distillation-based methods.
@@
DLM Acceleration. Most existing acceleration methods for masked dLLMs are either system level, such as enabling kv-cache and employing token dropping (Ma et al., 2025; Liu et al., 2025a; Hu et al., 2025; Wu et al., 2025) or algorithmic level that designs better sampling strategies. Among the latter, there are two approaches: (i) training-free logit-based heuristic decoders (Wu et al., 2025; Israel et al., 2025; Wei et al., 2025; Ben-Hamu et al., 2025) or (ii) distillation-based approaches that fit the off-policy trajectories produced by a base model (Chen et al., 2025; Qian et al., 2025). Heuristic samplers often rely on confidence, entropy or other related uncertainty signals computed from the logits, and have recently enabled substantial inference speedups in dLLMs via parallel decoding and KV caching (Wu et al., 2025; Hu et al., 2025). Distillation-based methods, such as pseudo-trajectory distillation in d3LLM (Qian et al., 2025), learnable parallel decoding in dParallel (Chen et al., 2025), and distribution-matching distillation in DiDi-Instruct (Zheng et al., 2025), can further increase decoding efficiency by finetuning the base model to sample multiple steps of the base model in one step, which essentially filters the modes of the base model. These methods depend on the quality and coverage of the teacher trajectories, which also may become off-policy as training progresses. Our approach explicitly addresses this mismatch by using on-policy rollouts with GRPO, while also encouraging mode seeking behavior.
@@
Denoising Trajectory Planning. Beyond heuristic confidence-based planning, recent works have explored learned planning modules for MDLM trajectories. These approaches differ fundamentally in their planning strategies. P2 (Peng et al., 2025) introduces explicit planning and denoising sub-stages, enabling iterative refinement of existing unmasked tokens. DCOLT (Huang et al., 2025) employs the Plackett-Luce model to generate sequential unmasking orders optimized for accuracy rather than parallel decoding efficiency. DiFFPO (Zhao et al., 2025a) learns adaptive per-prompt thresholds via off-policy RL but still selects tokens based on confidence heuristics. In the concurrent work, Jazbec et al. (2025) learns an unmasking policy head with GRPO on token confidences from logits.
@@ 6 Conclusion
We present dUltra, a framework for learning optimal unmasking strategies to accelerate masked diffusion language models through reinforcement learning. Motivated by the need for mode filtering and maximizing conditionally independent tokens per step, we train an unmasking planner head jointly with the base model using GRPO with on-policy rollouts and distillation rewards. Empirically, dUltra outperforms distillation baselines (d3LLM, dParallel) in both accuracy and NFEs, validating the superiority of our method. We found that learned unmasking order adapts to the task structure naturally: code generation exploits more parallelism locally by learning an unmasking order resembling the random masking strategy while math reasoning decodes multiple consecutive tokens sequentially and in parallel, exhibiting mode filtering behaviors. This work demonstrates that learned unmasking policies guided by mode filtering and conditional independence principles enable progress towards "diffusion supremacy" over autoregressive LLMs.
@@ A.1.1 Model Architecture
Our model extends the LLaDA-Instruct model [Nie et al., 2025] with an unmasking planner head. The architecture consists of:
Base Model. We use LLaDA [Nie et al., 2025], a masked diffusion language model with Llama-like architecture adapted for discrete diffusion. The model uses a special mask token and employs time-agnostic parametrization following recent findings that the optimal ELBO solution is independent of the time variable [Ou et al., 2025, Sahoo et al., 2024].
Unmasking Planner Head. Following the architecture in Huang et al. [2025], we introduce an unmasking planner head that takes the last-layer hidden states and outputs per-token unmasking probabilities:
@@ Appendix B Training Algorithm
Appendix C Quality-speed trade-off with multiplicative factors
The unmasking probability predicted by the unmasking planner head directly controls the expected number of tokens unmasked at each denoising step. This allows us to strike a balance between sampling efficiency (NFE) and task performance (accuracy). By multiplying the unmasking probability by a factor $\alpha$  [^4] For Table 1 and Table 2, we use the unscaled unmasking probability, i.e. $\alpha=1$., we are able to adjust the sampling speed at inference time without retraining the model. Figure˜9 illustrates the accuracy vs. NFE trade-off on all tasks for both block size 32 and block size 128.
Figure 10: Pareto frontier of dUltra-math vs dUltra-coding on math reasoning tasks.